# AudioToolAgent: An Agentic Framework for Audio-Language Models

### v1

@@ 4.1.2 Tools
-0.050.000.050.100.15Tavily Web Search [33]DuckDuckGo Web Search [34]StepAudio 2 [10]Parakeet TDT v3 [35]Granite Speech [36]Speaker Diarization [37]Phi-4 Multimodal [38]Gemma 3n [39]GPT4o-mini [21]DeSTA2.5 [9]Voxtral [12]Whisper v3 Turbo [13]Kimi Audio [8]AudSemThinker [40]Qwen2Audio [7]Gemini 2.5 Flash [5]AudioFlamingo 3 [11]Qwen2.5 Omni [6]Shapley valueAudio UnderstandingSpeech-to-TextSpeaker DiarizationAuxiliary
@@
Fig. 4: Shapley values per feature with standard error bars computed over 374 runs. Bars indicate the contribution of the audio-language model or API as tool to the overall system performance. Error bars are black.
@@
Qwen2.5 Omni [6] provides the highest contribution, followed by Audio Flamingo 3 [11] and Gemini 2.5 Flash [5]. The top-performing tools were incorporated in the configurations of this work (see Section 3.1) with two exceptions: Qwen2Audio [7] processes only 30 seconds of audio and audios in benchmarks are often longer, and AudSemThinker [40] shares its architecture with Qwen2.5 Omni.
4.2 Discussion and Future Work
AudioToolAgent’s performance depends on its underlying audio tools, creating both opportunities and challenges for future development. When tools produce inaccurate outputs, the agent may propagate these errors. To mitigate this, the agent uses a cross-validation approach, verifying information across multiple tools. By comparing answers to direct questions, the agent reduces reliance on single tool responses. Future work should explore advanced consensus mechanisms and uncertainty quantification to improve robustness against tool errors.
A practical limitation of AudioToolAgent is speed. Using multiple separate tools creates longer processing times than a single audio model. The agent calls tools sequentially and waits for each result, adding overhead. Running AudioToolAgent on separate machines mitigates this issue. Future research should focus on training agents to select optimal tool subsets for specific tasks, improving efficiency.
@@ 4.2 Discussion and Future Work
We tested the framework with web search integration using both DuckDuckGo API and Tavily’s proprietary search API (see Figure 4). Ablation studies revealed no consistent improvements, likely because the benchmarks focus on audio content and general information. These benchmarks mostly require historical information and common knowledge facts. Nevertheless, web search integration remains promising for real-world applications where external knowledge retrieval enhances performance. The tests also included the ability to extract parts of the audio and use them as input to the tools, but this did not improve performance. Future work should explore this approach. Expanding the tool ecosystem to include audio retrieval, audio generation, and analysis tools offers another valuable research direction.
Our evaluation approach uses existing benchmark results from model authors and benchmark maintainers (see Table 1) rather than reproducing all baseline numbers independently. This decision was made based on practical factors: high inference costs, API availability limitations, and the need for model-specific optimization to achieve peak performance. The unverified entries are marked where applicable.
5 Conclusion
@@
This paper introduced AudioToolAgent, a framework for multimodal audio understanding and reasoning where a central agent coordinates audio-language models as tools. The combination of GPT-4o, Gemini 2.5 Flash, Voxtral, Qwen2.5 Omni and Audio Flamingo 3 orchestrated by GPT-5 outperforms prior models on the MMAU, MMAR and MMAR-Pro benchmarks.
This framework establishes a new paradigm that combines the strengths of ALMs and LLMs. The ablation studies identified Qwen2.5 Omni and AudioFlamingo 3 as the most effective audio tools. Among LLMs, DeepSeek V3.1 and Kimi K2 demonstrated superior performance as orchestrating agents. This hybrid approach combines the audio processing of ALMs with the reasoning strengths of LLMs, creating a more flexible and powerful system than either model type could achieve alone.
This work opens several promising research directions. Future work includes expanding to other tasks and developing learned tool selection policies, improving the speed of the framework, and exploring the use of web search and other analysis tools.
6 Acknowledgment
This work was supported by the Dutch Research Council (NWO 406.20.GO.030 to Prof. Elia Formisano), the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-14224, Data Science Research Infrastructure (DSRI; Maastricht University) and the Dutch Province of Limburg.

### v2

@@ IV-A2 Tools
-0.050.000.050.100.15Tavily Web Search [32]DuckDuckGo Web Search [9]Parakeet TDT v3 [6]Granite Speech [30]Speaker Diarization [4]Phi-4 Multimodal [19]Gemma 3n [29]GPT4o-mini [21]DeSTA2.5 [16]Voxtral [15]Whisper v3 Turbo [26]Kimi Audio [12]AudSemThinker ([38])Qwen2Audio [5]Qwen3 Omni [43]MiMo Audio [35]Gemini 2.5 Flash [7]AudioFlamingo 3 [11]Qwen2.5 Omni [42]Gemini 3 Pro [10]Shapley valueAudio UnderstandingSpeech-to-TextSpeaker DiarizationAuxiliary
@@
Figure 4: Shapley values per feature with standard error bars computed over 490 sampled configurations. Bars indicate the contribution of each audio tool to the overall system performance. Error bars are black.
@@
The best performing tools were incorporated into the configurations of this work (see Section III-A), with two exceptions: Qwen2Audio [5] processes only 30 seconds of audio and the audios in the benchmarks are often longer, and AudSemThinker [38] shares its architecture with Qwen2.5 Omni.
[…]
IV-B Discussion and Future Work
The multi-hop consensus mechanism mitigates tool errors by querying multiple models and issuing targeted follow-up queries when outputs conflict, producing higher agreement rates than any single model. This approach implements a form of consensus aggregation where majority agreement across independent models provides a more reliable signal than any single model. Individual models may exhibit biases or hallucinations, but combining evidence from multiple sources reduces the impact of such errors.
The agent ablation reveals that tool-calling proficiency varies across models: the best closed-source agent (Opus 4.5: 0.752) outperforms the best open-source agent (Qwen 3 235B: 0.710) by 4.2 percentage points, while the gap between the weakest and strongest agent spans 13.6 percentage points. Orchestration quality depends on training methodology rather than model scale alone.
[…]
A practical limitation is latency. Individual audio-language models process audio in 1–5 seconds per example, depending on model size. In contrast, AudioToolAgent requires 30 seconds to one minute per example because the framework coordinates multiple models sequentially, with the agent waiting for each tool’s response before proceeding. Four strategies mitigate this overhead. First, the orchestrator can invoke multiple tool calls simultaneously when the tools are independent, reducing the number of sequential round-trips. Second, concurrent batching across examples allows the orchestrator to process multiple audio questions in parallel: while waiting for tool responses on one example, the orchestrator can initiate tool calls for other examples. Third, deploying multiple replicas of tools across GPUs enables load balancing and reduces queuing delays. Fourth, using API endpoints from external providers eliminates the need to run local GPU infrastructure while enabling parallel API calls at low cost per query. Because no training is involved, the total compute cost remains modest compared to fine-tuning or training new models. Future research should focus on developing learned tool selection policies and lower-latency orchestration strategies to reduce sequential dependencies.
@@ IV-B Discussion and Future Work
The experiments tested the framework with web search integration using both DuckDuckGo API and Tavily’s proprietary search API (see Figure 4). Ablation studies did not reveal consistent improvements, likely because the benchmarks focus on audio content requiring historical information and common knowledge facts. Web search integration may benefit real-world tasks where external knowledge retrieval is required beyond the audio content.
A next step is combining native multimodal processing with tool-calling capabilities, merging the efficiency of direct audio understanding with the extensibility of frameworks like AudioToolAgent.
V Conclusion
@@
This paper introduced AudioToolAgent, a framework for multimodal audio understanding and reasoning where a central agent coordinates audio-language models as tools. The combination of Gemini 3 Pro, Qwen2.5 Omni, Audio Flamingo 3, and Whisper orchestrated by Gemini 3 Pro outperforms prior models on the MMAU, MMAR and MMAU-Pro benchmarks.
This framework combines the strengths of audio-language models and large language models. Ablation studies identified Gemini 3 Pro, Qwen2.5 Omni, and AudioFlamingo 3 as the highest-contributing audio tools across 20 evaluated tools. Among 12 LLMs tested as orchestrating agents, Claude Opus 4.5 and Gemini 3 Pro demonstrated the strongest performance, while open-source models like Qwen 3 235B achieved competitive results. This hybrid approach combines the audio processing of ALMs with the reasoning strengths of LLMs, creating a modular system in which tools and agents can be swapped independently.
The modular design enables integration of new tools and agents as they become available, without retraining. Future work includes developing learned tool selection policies, improving orchestration speed, and expanding the tool ecosystem to include audio retrieval and generation tools.
VI Acknowledgments
This work was supported by the Dutch Research Council (NWO 406.20.GO.030 to Prof. Elia Formisano), the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-14224. GW and JJ acknowledge funding by the Federal Ministry of Education and Research of Germany (BMBF) under grant no. 01IS24085C (OPENHAFM), under the grant 16HPC117K (MINERVA) and under the grant no. 01IS22094B (WestAI - AI Service Center West), as well as funding by EU Horizon under grant no. 101214398 (ELLIOT) and co-funding by EU from EuroHPC Joint Undertaking programm under grant no. 101182737 (MINERVA) and from Digital Europe Programme under grant no. 101195233 (openEuroLLM).